# Nonparametric Reaction Coordinate Optimization with Histories: A Framework for Rare Event Dynamics

### v1

@@ Abstract
Rare but critical events in complex systems, such as protein folding, chemical reactions, disease progression, and extreme weather or climate phenomena, are governed by complex, high-dimensional, stochastic dynamics. Identifying an optimal reaction coordinate (RC) that accurately captures the progress of these dynamics is crucial for understanding and simulating such processes. This work introduces a nonparametric RC optimization framework that incorporates trajectory histories, enabling robust analysis even for irregular or incomplete data. The power of the method is demonstrated through increasingly challenging analyses of protein folding dynamics, where it provides accurate committor estimates that pass a stringent validation test and yield high-resolution free energy profiles. Its generality is further illustrated through applications to dynamics in phase space, a conceptual ocean circulation model, and a longitudinal clinical dataset. These results demonstrate that rare event dynamics can be accurately characterized without exhaustive sampling of the configuration space, establishing a general, flexible, and robust framework for analyzing complex dynamical systems and longitudinal datasets.
@@
Many rare yet important events are governed by complex, high-dimensional stochastic dynamics. Understanding and simulating such events can be greatly facilitated by identifying a variable or index, often called a reaction coordinate (RC), that accurately captures the progress of the system’s dynamics [1, 2]. In chemical kinetics, for instance, understanding rare events is important for such processes as chemical reactions, crystallization, and protein folding and aggregation. In these systems, knowing an appropriate RC enables the identification of transition states — the critical points or bottlenecks that control the reaction dynamics. Moreover, an accurate RC can be used to accelerate or enhance the sampling of rare events in simulations [3, 4].
@@
For an RC to provide a faithful description of kinetics and to be practically useful for sampling, it should be optimally selected to minimize information loss resulting from dimensionality reduction. One prominent example of such an optimal RC is the committor function (also known as the splitting probability, or $p_{\text{fold}}$), which quantifies the likelihood that a given configuration will evolve towards a particular final state before reaching an alternative state. While originally developed in the context of chemical kinetics [5, 6, 7], the concept of the committor has found broad applicability across diverse fields. For example, it can be used to assess the probability of extreme weather or climate events [8, 9], serve as an optimal biomarker to predict a patient’s likelihood of recovery based on current conditions [10], or estimate the probability of winning a game of chess from a given position [11].
@@
Finding the committor for realistic systems is a notoriously difficult problem. The committor is a complex, non-linear function, formally defined as the solution of a high-dimensional partial differential equation over the system’s configuration space. Many promising approaches have been proposed to determine the committor, with recent advances driven by machine learning (ML) [12, 13, 14, 15, 16, 17, 18, 19, 20, 19, 8, 21, 3, 4, 22, 23, 24]. Despite the remarkable success of ML techniques in a variety of domains, for example, protein structure prediction [25] or design [26], progress in applying ML to learn the committor in complex systems has so far been relatively limited. Most existing studies have focused either on low-dimensional model systems or on realistic systems with relatively simple RCs and free energy landscapes featuring a single barrier. Reproducible analyses of systems with complex RCs and free energy landscapes involving multiple metastable states, which can serve as rigorous benchmarks to drive the field’s development, remain scarce [27]. These observations highlight the unique methodological challenges in determining the committor and analyzing rare event dynamics, suggesting that alternative approaches beyond standard ML techniques may be necessary [28, 29].
An additional complication arises from the nature of real-world data. While computer-simulated trajectories are typically regular, many real-world datasets involve highly irregular trajectories, often with an incomplete set of variables and missing data. Such challenges are common across domains. Clinical datasets are often incomplete and complicated by patient dropouts. Weather observations combine heterogeneous sources with sparse, uneven sampling [30]. Single-molecule experiments may provide only very low-dimensional projections of the underlying dynamics [31, 32, 33]. These irregularities make the determination of the committor function particularly difficult.
Here, we present an approach for robust and accurate determination of the committor function (and other optimal RCs) from irregular and incomplete longitudinal datasets. The approach yields reproducible results in the analyses of a realistic protein folding trajectory, characterized by complex RCs and corresponding free energy landscapes. It is further illustrated on diverse systems, including dynamics in phase space, conceptual ocean circulation models, and a clinical dataset.
@@ Results
To introduce the idea of RC optimization, consider stochastic dynamics in a multidimensional configuration space $\vec{X}$, describing a reaction between two specified boundary states, $A$ and $B$. Introduce an RC, $r(\vec{X}):\vec{X}\rightarrow[0,1]$, which projects $\vec{X}$ onto a one-dimensional coordinate while satisfying boundary conditions $r(\vec{X}\in A)=0$ and $r(\vec{X}\in B)=1$. Dynamics projected onto an RC are generally non-Markovian due to the loss of information associated with the dimensionality reduction. This non-Markovian behavior can be accurately described using the generalized Langevin equation (GLE) with a memory kernel, which can be derived via the Mori-Zwanzig formalism [34, 35]. In principle, the GLE can be derived for any reasonable coordinate; however, computing the memory kernel in practice is very difficult [36, 37].
@@
We consider two scenarios for describing the dynamics, each defined by a different set of boundary conditions. In the first scenario (Figs. 7a–b), each hospital admission is treated as an independent trajectory. Surrogate endpoints are defined as follows: if a patient experiences AKI3, identified using a standard clinical algorithm [60], the trajectory ends in state $B$; otherwise, it ends in state $A$. This design ensures that every trajectory terminates in a boundary state, allowing accurate estimation of observed probabilities. The optimal RC in this scenario quantifies the likelihood of developing AKI3 during a hospital stay. Figure 7a shows that $Z_{q}$ remains relatively constant, except in the region of very small $q$. This indicates that the computed RC closely approximates the committor function, apart from that low-$q$ region. Predicted and observed probabilities (Fig. 7b) are in strong agreement. In contrast, optimization without histories yields significantly worse results (Supplementary Fig. LABEL:fig:S5).
@@ Discussion
A key strength of the proposed framework is its generality. Unlike many alternative methods, it does not rely on common assumptions such as detailed balance, equilibrium sampling, or overdamped dynamics. It is applicable to systems with missing variables, irregular time series, or incomplete observables, and to cases where exhaustive sampling is infeasible. In particular, it can be used to analyze systems governed by Fokker–Planck equations with position-dependent diffusion coefficients, underdamped Langevin dynamics, or rare event dynamics in exponentially large configuration spaces.
@@
In comparison with GLE approaches, both the proposed approach and GLE approaches incorporate memory effects, but in fundamentally different ways. The proposed approach leverages memory, or more precisely the system’s history, to determine a complex, nonlinear mapping from the history space to a position along the optimal RC. Once the position is determined, the dynamics are modeled using a simple diffusive model, yielding a clear and intuitive description of the system’s overall dynamics. In contrast, GLE-based approaches employ a simple mapping to the RC while embedding all the dynamics complexity within the memory kernel. This results in non-Markovian dynamics that depend on the entire history of systems evolution. While the memory kernel formally encapsulates all relevant dynamical information, it often acts as a “black box,” making mechanistic interpretation, such as identifying transition states, more challenging without explicit simulation.

### v2

# arXiv:2508.07326v2[physics.chem-ph] 03 Mar 2026

@@ Abstract
Rare but critical events in complex systems, such as protein folding, chemical reactions, disease progression, and extreme weather or climate phenomena, are governed by complex, high-dimensional, stochastic dynamics. Identifying an optimal reaction coordinate (RC) that accurately captures the progress of these dynamics is crucial for understanding and simulating such processes. However, determining an optimal RC for realistic systems is notoriously difficult, due to methodological challenges that limit the success of standard machine learning techniques. These challenges include the absence of ground truth, the lack of a loss function for general nonequilibrium dynamics, the difficulty of selecting expressive neural network architectures that avoid overfitting, the irregular and incomplete nature of many real world trajectories, limited sampling and the extreme data imbalance inherent in rare event problems. Here, we introduce a nonparametric RC optimization framework that incorporates trajectory histories and circumvents these challenges, enabling robust analysis of irregular or incomplete data without requiring extensive sampling. The power of the method is demonstrated through increasingly challenging analyses of protein folding dynamics, where it yields accurate committor estimates that pass stringent validation tests and produce high resolution free energy profiles. Its generality is further illustrated through applications to phase space dynamics, a conceptual ocean circulation model, and a longitudinal clinical dataset. These results demonstrate that rare event dynamics can be accurately characterized without extensive sampling of the configuration space, establishing a general, flexible, and robust framework for analyzing complex dynamical systems and longitudinal datasets.
@@
Many rare yet important events are governed by complex, high-dimensional stochastic dynamics. Understanding and simulating such events can be greatly simplified by identifying a variable or index, often called a reaction coordinate (RC), that accurately captures the progress of the system’s dynamics [1, 2]. In chemical kinetics, for instance, understanding rare events is important for such processes as chemical reactions, crystallization, and protein folding and aggregation. In these systems, knowing an appropriate RC enables the identification of transition states, the critical points or bottlenecks, that control the reaction dynamics. Moreover, an accurate RC can be used to accelerate or enhance the sampling of rare events in simulations [3, 4].
@@
For an RC to provide a faithful description of kinetics and to be practically useful for sampling, it should be optimally selected to minimize information loss resulting from dimensionality reduction. One prominent example of such an optimal RC is the committor function (also known as the splitting probability, or $p_{\text{fold}}$), which quantifies the likelihood that a given configuration will evolve towards a particular final state before reaching an alternative state. While originally developed in the context of chemical kinetics [5, 6, 7], the concept of the committor has since been applied to problems across other fields. For example, it has been used to assess the probability of extreme weather or climate events [8, 9], to serve as an optimal biomarker to predict a patient’s likelihood of recovery based on current conditions [10], and to estimate the probability of winning a game of chess from a given position [11].
@@
Finding the committor for realistic systems is a notoriously difficult problem. The committor is a complex, non-linear function, formally defined as the solution of a high-dimensional partial differential equation over the system’s configuration space. Many promising approaches have been proposed to determine the committor, with recent advances driven by machine learning (ML) [12, 13, 14, 15, 16, 17, 18, 19, 20, 19, 8, 21, 3, 4, 22, 23, 24]. Despite the remarkable success of ML techniques in a variety of domains, for example, protein structure prediction [25] or design [26], progress in applying ML to learn the committor in complex systems has so far been relatively limited. Most existing studies have focused either on low-dimensional model systems or on realistic systems with relatively simple RCs and free energy landscapes featuring a single barrier. Reproducible analyses of systems with complex RCs and free energy landscapes involving multiple metastable states, which can serve as rigorous benchmarks to drive the field’s development, remain scarce [27]. This limited progress reflects the fact that determining the committor and analysing rare-event dynamics present a set of unique methodological challenges that make the straightforward application of standard ML techniques difficult. Next, we briefly highlight the main challenges that complicate committor estimation and explain how they fundamentally differentiate this problem from generic ML tasks.
No ground truth. For complex, realistic systems, the true values of the committor function are generally unknown, which makes it difficult to assess the accuracy of computed committors. In principle, one can evaluate the committor directly from its definition by initiating a large number of trajectories from a given configuration and measuring the fraction that reach state B before state A [6, 1] . However, this brute-force approach is prohibitively expensive and can only be applied to a limited number of configurations, not to entire trajectories, and is infeasible in applications where generating new trajectories is impossible, such as analyses of longitudinal patient data. This limitation likely explains why only a small number of studies have attempted to quantify the accuracy of computed committors in complex, realistic systems.
No valid loss/objective function for general nonequilibrium dynamics. While there are valid loss functions for certain specific cases, where the committor is the exact minimizer, such as equilibrium dynamics [7, 28], Langevin dynamics with a constant diffusion coefficient [4, 29], or nonequilibrium settings in which every trajectory reaches a boundary state [22], no such loss function is known for the general case of non-equilibrium dynamics, e.g., ensembles of short trajectories, often encountered in practical applications. As a result, standard ML practices such as train/test splits or cross-validation used to evaluate model performance or detect overfitting cannot be directly applied.
NN architecture. Selecting an appropriate neural network architecture is challenging: on the one hand, the network must be sufficiently expressive to approximate the highly complex, high-dimensional committor; on the other hand, it must contain a manageable number of parameters to reduce the risk of overfitting [28]. This difficulty is compounded by the fact that it is hard to assess approximation accuracy or reliably detect overfitting in the absence of a true loss function or ground truth.
Irregular data. Empirical trajectories, in contrast to those generated in controlled computer simulations, are usually highly irregular. For example, trajectories may be interval censored or right censored when patients miss appointments or drop out of a study. They often contain missing entries, inconsistent sampling intervals, or corrupted measurements. Weather datasets exhibit similar challenges: observations are collected from heterogeneous sources with sparse, uneven sampling [30]. Single-molecule experiments may provide only very low-dimensional projections of the underlying dynamics [31, 32, 33]. Standard ML methods are generally sensitive to such imperfections in the input data, making accurate committor estimation considerably more challenging in these settings.
Rare events. Rare events dynamics datasets are, by definition, strongly imbalanced; the configurations associated with the events of interest, such as folding-unfolding transitions in protein folding trajectories, constitute only a tiny fraction of the full dataset. Consequently, performance must be evaluated using metrics that are sensitive to the accuracy of the committor specifically in these rare regions, since their contribution to aggregate metrics (e.g., mean squared error) may be negligible. For example, if only 1% of configurations are misclassified, this may appear acceptable; however, for realistic ratios, e.g., when the population of the transition state (TS) is 1/1000 that of the native basin, this would imply that ten out of eleven putative TS configurations are actually native basin states misclassified as TS.
A second challenge is achieving convergence during batch based optimization, a standard ML technique [34, 28]. Most batches will contain no or at most a few rare events of interest, making the gradient estimate uninformative.
Finally, in ML one typically requires that the model (the learned committor) generalize to unseen data. However, generalizability requires that the train and test data follow the same underlying distribution, which in turn requires extensive and representative sampling. Ensuring such coverage is particularly difficult for rare events in high-dimensional configuration spaces [35].
Progress in this research area requires addressing all of these challenges. We propose the framework of nonparametric RC optimization with histories as one possible way to do so. Specifically, we present methods for the robust and accurate determination of the committor function (and other optimal RCs) from irregular and incomplete longitudinal datasets. The non-parametric nature of the framework allows the committor to be approximated arbitrarily closely, and using the entire trajectory at each iteration avoids the small batch issues for rare events. A key ingredient of the framework is the set of highly sensitive validation criteria that assesses the accuracy of the putative committor (and other optimal RCs). Importantly, the criteria are not based on a single aggregate quantity: they verify that the putative RC satisfies the corresponding equation for every value of the RC and across all relevant time scales. Because the validation does not rely on train/test splits, it does not assume extensive or representative sampling of the configuration space. This, in turn, makes it possible to separate two tasks: (i) accurately determining the committor within the sampled region, and (ii) learning a global committor that generalizes to unseen configurations or the entire configuration space. The former is sufficient for simulation analysis and enhanced sampling [36], whereas the latter would require exponentially more data and is typically infeasible.
The power of the framework is thoroughly evaluated through increasingly challenging analyses of realistic protein folding dynamics, where it yields accurate committor estimates that pass the stringent validation criterion and produce high-resolution free energy profiles. Its generality is further illustrated across diverse systems, including phase space dynamics, conceptual ocean circulation models, and a clinical longitudinal dataset.
@@ Results
To introduce the idea of RC optimization, consider stocshastic dynamics in a multidimensional configuration space $\vec{X}$, describing a reaction between two specified boundary states, $A$ and $B$. Introduce an RC, $r(\vec{X}):\vec{X}\rightarrow[0,1]$, which projects $\vec{X}$ onto a one-dimensional coordinate while satisfying boundary conditions $r(\vec{X}\in A)=0$ and $r(\vec{X}\in B)=1$. Dynamics projected onto an RC are generally non-Markovian due to the loss of information associated with the dimensionality reduction. This non-Markovian behavior can be accurately described using the generalized Langevin equation (GLE) with a memory kernel, which can be derived via the Mori-Zwanzig formalism [37, 38]. In principle, the GLE can be derived for any reasonable coordinate; however, computing the memory kernel in practice is very difficult [39, 40].
@@
We consider two scenarios for describing the dynamics, each defined by a different set of boundary conditions. In the first scenario (Figs. 7a-b), each hospital admission is treated as an independent trajectory. Surrogate endpoints are defined as follows: if a patient experiences AKI3, identified using a standard clinical algorithm [63], the trajectory ends in state $B$; otherwise, it ends in state $A$. This design ensures that every trajectory terminates in a boundary state, allowing accurate estimation of observed probabilities. The optimal RC in this scenario quantifies the likelihood of developing AKI3 during a hospital stay. Figure 7a shows that $Z_{q}$ remains relatively constant, except in the region of very small $q$. This indicates that the computed RC closely approximates the committor function, apart from that low-$q$ region. Predicted and observed probabilities (Fig. 7b) are in strong agreement. In contrast, optimization without histories yields significantly worse results (Supplementary Fig. LABEL:fig:S5).
@@ Discussion
A key strength of the proposed framework is its generality. Unlike many alternative methods, it does not rely on common assumptions such as detailed balance, equilibrium sampling, or overdamped dynamics. It is applicable to systems with missing variables, irregular time series, or incomplete observables, and to cases where exhaustive sampling is infeasible. In particular, it can be used to analyze systems governed by Fokker-Planck equations with position-dependent diffusion coefficients, underdamped Langevin dynamics, or rare event dynamics in exponentially large configuration spaces.
@@
In comparison with GLE approaches, both the proposed approach and GLE approaches incorporate memory effects, but in fundamentally different ways. The proposed approach leverages memory, or more precisely the system’s history, to determine a complex, nonlinear mapping from the history space to a position along the optimal RC. Once the position is determined, the dynamics are modeled using a simple diffusive model, yielding a clear and intuitive description of the system’s overall dynamics. In contrast, GLE-based approaches employ a simple mapping to the RC while embedding all the dynamics complexity within the memory kernel. This results in non-Markovian dynamics that depend on the entire history of systems evolution. While the memory kernel formally encapsulates all relevant dynamical information, it often acts as a “black box”, making mechanistic interpretation, such as identifying transition states, more challenging without explicit simulation.